OpenStudy (adamaero):

Medal - What's magnitude of just an imaginary number with no real part? I mean, I have iωL in a denominator...

OpenStudy (kainui):

The magnitude of every imaginary number z is given by the square root of the number multiplied with it's complex conjugate. \[|z| = \sqrt{zz^*}\] It's clear to see why this is true cause of the pythagorean theorem: \[z= a+bi\]\[z=a-bi\] \[zz^* = (a+bi)(a-bi) = a^2+abi-abi+(i)(-i)b^2 = a^2 + b^2\] Hopefully that helps

OpenStudy (adamaero):

Am I supposed to take it out when finding magnitude? e.g., ||1/(iωL)|| =?= 1/(ωL)

OpenStudy (adamaero):

Am I missing something conceptually? All I think is i = \[\sqrt{-1}\]

OpenStudy (kainui):

You don't take it out, just follow the process, multiplying a number by its complex conjugate always gives a real number. Try it out

OpenStudy (kainui):

If you don't like that cause i is in the denominator then I can teach you a trick to juggle it up into the numerator: Since \(\sqrt{-1} = i\) we can just square it to get -1, nothing crazy: \[i*i =-1\] Now multiply both sides by -1 \[-i*i=1\] Divide both sides by i \[-i = \frac{1}{i}\] So any time you see \(\frac{1}{i}\) replace it with \(-i\) since they're equal.

OpenStudy (adamaero):

but I can't have imaginary numbers as a part a magnitude